OpenStudy (anonymous):

The equation sin ( (PI/3) - x ) is equal to ____. A. 1/2( sqrt(3) cosx + sinx) B. 1/2(sqrt(3) cosx - sinx) C. 1/2(cosx- sqrt(3) sinx) D. 1/2 (cosx + sqrt(3) sinx)

OpenStudy (anonymous):

use sin(A-B) = sinA*cosB - sinB*cosA

OpenStudy (anonymous):

can you show me how to do it? if you have time of course..... i don't know how to do it

OpenStudy (anonymous):

Use the trig identity: sin (a - b) = sin a*cos b - sin b*cos a sin Pi/3 = V3/2 ; cos Pi/3 = 1/2 sin(Pi/3 - x) = (V3/2)cos x - (sin x)/2 = 1/2(V3*cos x - sin x)

OpenStudy (anonymous):

i see what you did; thank you so much!